OpenStudy (anonymous):

A coin bank contains $17.00 in pennies and nickels. If there are 1140 coins in the bank, how many coins are nickels? a. 134 b. 137 c. 140 d. 143

OpenStudy (igreen):

We can write a system of equations. x + y = 1140 0.01x + 0.05y = 17

OpenStudy (bloomlocke367):

okay, what you can do is set up a system of equations... \[0.01p+0.05n=17\]\[p+n=1140\] p=pennies, n=nickels

OpenStudy (bloomlocke367):

gah, iGreen, stop stealing my thunder XD lol jk

OpenStudy (igreen):

We can solve using substitution. x + y = 1140 Subtract 'y' to both sides: x = -y + 1140 Plug in '-y + 1140' for 'x' in the 2nd equation: 0.01(-y + 1140) + 0.05y = 17 Distribute 0.01 into the parenthesis: -0.01y + 11.4 + 0.05y = 17 Combine like terms: 0.04y + 11.4 = 17 Subtract 11.4 to both sides, what's 17 - 11.4?

OpenStudy (anonymous):

5.6

OpenStudy (bloomlocke367):

then divide by 0.04

OpenStudy (anonymous):

so its c right

OpenStudy (anonymous):

do you still need help @~cuddle_monster~

OpenStudy (anonymous):

i dont think so they help alot but thanks for the offer

OpenStudy (igreen):

Yeah, 5.6 / 0.04 = ?

OpenStudy (igreen):

Yep, C is correct.. 5.6 / 0.04 = 140

OpenStudy (anonymous):

thanks you guys

OpenStudy (bloomlocke367):

np